OpenStudy (anonymous):

If f (x) = |x-5| + 3, find f (2) (points 1) A. 11 B. 6 C. 4 D. 0 Help? Please? :-)

Nnesha (nnesha):

replace x by 2

Nnesha (nnesha):

f(2) mean that x = 2 so then u can substitute into x

OpenStudy (anonymous):

If you need more help just mention my name :)

OpenStudy (usukidoll):

by absolute value rules we either have f(x) = x-5+3 f(x) = x-2 f(2) = 0 or f(x) = -x+5+3 f(x) = -x+8 f(2) = 6

OpenStudy (usukidoll):

or.. we can just directly subsitute it and take the absolute value f(x) = [x-5]+3 f(2) = [2-5]+3 f(2) = [-3]+3 f(2) = 3+3 =6

OpenStudy (anonymous):

Thank you guys for all your help :)